What point is the intersection of the graphs of x2 - y2 = 5 and 2x2 + y2 = 22?

Answer by Alan3354(69443)

What point is the intersection of the graphs of x2 - y2 = 5 and 2x2 + y2 = 22?
-----------
There are 4 of them.
x2 - y2 = 5 --> x^2 = y^2 + 5
Sub for y^2 in the other equation and solve for x.
It's tedious, but you need the practice more than I do.
email your work via the thank you note to check it, or for help.